# A deep learning model for chemical shieldings in molecular organic solids including anisotropy

### v2

@@ Acknowledgements.
The ensemble uncertainty estimates we obtain are miscalibrated (as it is often the case [88, 89] in uncertainty quantification schemes for neural networks), i.e. the uncertainty estimates across all samples of a hold-out validation or test set model uncertainty estimates are systematically over- or underconfident. It is possible to rescale the uncertainty estimates, and correct the miscalibration by computing a global, input-independent scaling factor [57, 96] (an empirical correction that is similar in spirit to temperature scaling[88] a classical method in uncertainty quantification). The scaling factor can be computed via a closed form expression, that can be derived by maximizing the Gaussian negative log likelihood on a hold out test set given predicted uncertainties and squared prediction errors $z^{2}_{i}(A)$.
$\alpha^{2}=\sum^{N_{\text{samples}}}_{i=1}\frac{z^{2}_{i}(A)}{\sigma^{2}_{i,\text{pred}}(A)}$ (2)
The final, calibrated, uncertainties are then obtained as $\alpha^{2}\sigma^{2}_{\text{pred}}(A)$ although in practice we scale the ensemble members around their mean, which simplifies uncertainty propagation. [86] We find that for isotropic chemical shieldings and shielding tensor principal components the calibration factors $\alpha$ are moderate and always close to 111, indicating that the ShiftML3 ensemble uncertainty estimates, are already well calibrated.
We evaluate the uncertainty estimates of the ShiftML3 model on the CSD-test set and plot the predicted uncertainties of 13C shieldings vs the actual prediction errors. In Fig. S3 we show parity plots of predicted and calibrated uncertainties of the ShiftML3 model and the actual errors on a hold out test set.
Figure S3: Predicted uncertainties of 13C isotropic chemical shieldings $\text{std}_{\text{pred}}[\sigma_{\text{iso}}]$ vs the actual prediction errors $\text{z}[\sigma_{\text{iso}}]$ (a) and predicted uncertainties of the corresponding principal components $\text{std}_{\text{pred}}[\sigma_{\text{PAS}}]$ vs the actual component prediction errors $\text{z}[\sigma_{\text{PAS}}]$ (b). The grey lines indicate quantiles of the expected error distribution for a given predicted uncertainty values and are intended to guide the readers eye: For example, on the test set, for well calibrated uncertainty estimates there should be 90% of the samples between the two outmost lines as they correspond to quantile lines of 95% and 5% of the folded log normal distribution.
@@
Only the symmetric part enters directly into the NMR spectrum, whilst the antisymmetric part effects relaxation times. [97] Therefore, in the main part of this manuscript we have focussed on evaluating and utilizing the CSA, $\sigma_{\text{PAS}}$ which are the principal components of $\sigma_{\text{sym}}$. ShiftML3 can predict the full chemical shielding tensor, containing symmetric and antisymmetric parts. In Figure S4 we show parity plots of the symmetric and antisymmetric components $\sigma_{ij,\text{sym}}$ and $\sigma_{ij,\text{antisym}}$ of the 13C and 17O chemical shielding tensors.
@@
In Table S4 we list the prediction accuracies of ShiftML on CSD-test separately on antisymmetric and symmetric components of the chemical shielding tensor.
@@
Figure S5: Parity plots of symmetric and antisymmetric chemical shielding tensor components $\sigma_{ij,\text{sym}}$, $\sigma_{ij,\text{antisym}}$) for 13C (a) and 17O (b) - combined parity plots.
@@
Figure S6: Parity plots of principal tensors components $\sigma_{\text{PAS}}$ and symmetric tensor components $\sigma_{ij}$ for 13C (a,b) and 15N (c,d)
In Figure S7 we show parity plots of the prinicpal tensor components resolved for individual principal components $\sigma_{11}$,$\sigma_{22}$ and $\sigma_{33}$. In Table S5 we list prediction accuracies of the individual principal components.
@@
Table S5: Predicted accuracy of $\sigma_{\text{PAS}}$ of (sML3) on the CSD-test hold-out set, resolved for individual principal components $\sigma_{\text{11}}$, $\sigma_{\text{22}}$ and $\sigma_{\text{33}}$.
Species MAE [ppm] RMSE [ppm] RMSE [%]
@@
For 13C and 15N, the accuracies were evaluated using the benchmark of 36 structures from Ramos et al. [14], with geometries optimized at the PBE-D3(BJ) level of theory [98]. For the carbon CSA parameters, we used the benchmark consisting of 19 structures from Hartman et al. [72].
Evaluation of 1H was done on the same 13 structures where ShiftML2 was evaluated [49] on the same geometry optimized structures. Predicted shieldings of methyl or primary amine groups were averaged out accounting for rapid rotation, resulting in the averaged sites have equivalent chemical shielding. Ambiguous experimental assignments were solved by iterating along all possible permutations of those unassigned shifts and taking the assignment with the lowest RMSE between predicted and experimental chemical shifts such as for methylene groups. Only the permutation with the lowest RMSE entered the global linear regression. We note that the reported RMSE comparing prediction to experiment for ShiftML2 is modestly different than the original work. This difference originates from the use of a global regression as compared to on-the-fly regression used in the original work. For 13C and 15N, the accuracy was were evaluated from the benchmark consisting of of 36 structures from Ramos et al. [14] on geometries optimized on the PBE-D3(BJ)level of theory [98]. For the carbon CSA parameters, we used the benchmark consisting of 19 structures from Hartman et al. [72]
Table S6: RMSE and Linear regression parameters are for the DFT, ShiftML2, and ShiftML3 prediction and experimental chemical shifts from the respective benchmark.
@@
A direct comparsion of the ShiftML3 prediction accuracies with ShiftML2 reveals that it is more accurate for predicting isotropic chemical shieldings across all relevant elements. To the best of our knowledge, no other machine learning model has been trained on the last iteration of the ShiftML datasets, including thermally distorted structures and additional chemical elements. Unzueta et al. trained GNN models [87] and Liu and coworkers a multiresolution 3D-DenseNet architecture on the CSD-2K dataset [46], an early iteration of the ShiftML training sets. An alternative model construction strategy is the NMRNet from Xu and coworkers, that finetuned a transformer model for the chemical shielding prediction in liquids on the CSD-2K set [99]. All three works report evaluation accuracies on the CSD-500 set, a set of DFT relaxed organic crystals, containing at most H,C,N and O. Unfortunately the number of shared structures between the CSD-test and the CSD-500 test set is too small (about 10 structures) to make direct comparisons between the accuracies of ShiftML3 and other neural network architectures. Shielding computations are highly sensitive to k-space discretization and choice of pseudopotentials [100]. CSD-2K/CSD-500 was computed with other DFT convergence parameters than the ShiftML2/ShiftML3 dataset, hampering severely a direct comparison between all models - we regularly find that shielding values from computations employing different, yet reasonable convergence parameters, differ (in terms of their RMSEs) nearly as much as the ShiftML3 predicted shieldings differ from the CSD test reference calculations. Instead, we evaluate ShiftML2 and ShiftML3 on 351 structures in our test set that only contain H,C,N,O and are not MD distorted structures (containing 7536 1H, 6546 13C, 665 15N and 1311 17O test environments) and compare their relative improvements in accuracies against the other models, to give some context to the gains in accuracy we achieve with ShiftML3. In Table S7, we compare prediction accuracies of the original ShiftML1.0 model, Unzuetas ensemble of GNNs, Liu’s MR-3D-DenseNet and compare them with the prediction accuracies of ShiftML2 and ShiftML3 on the H,C,N,O-relaxed subselection of the CSD-test set.
Table S7: Test set RMSEs (ppm) of the original ShiftML1.0 model (KRR [45]), Unzuetas GNNs (GNN [87]) and Liu’s MR-3D-DenseNet (DenseNet [46]) and Xu’s NMRNet (NMRNet [99]) on the CSD-500 test set. We compare these values with the prediction accuracies of ShiftML2 (sML2) and ShiftML3 (sML3) on a comparable subset of the CSD test set. Note that KRR, GNN, Densenet and NMRNet are evaluated on the CSD-500 test set. As discussed in the text, due to the small overlap of CSD-500 and the CSD test set, we choose to evaluate sML2 and sML3 on a subset of the CSD test set containing only H,C,N,O structures that are geometry relaxed. (*) Denotes, that the results of sML2 and sML3 were obtained on that subselection.

### v3

@@ Acknowledgements.
Ensemble uncertainty estimates are typically miscalibrated (as is often the case [88, 89] in uncertainty quantification schemes for neural networks), i.e. the uncertainty estimates across all samples of a hold-out validation or test set model uncertainty estimates are systematically over- or underconfident. It is possible to rescale the uncertainty estimates, and correct the miscalibration by computing a global, input-independent scaling factor [57, 96] (an empirical correction that is similar in spirit to temperature scaling[88] a classical method in uncertainty quantification). The scaling factor can be computed via a closed-form expression, that can be derived by maximizing the Gaussian negative log-likelihood on a hold-out test set given predicted uncertainties $\sigma^{2}_{\text{pred}}(A_{i})$ and squared prediction errors $z^{2}(A_{i})$ of some environment $A_{i}$.
$\alpha^{2}=\frac{1}{N_{\text{samples}}}\sum^{N_{\text{samples}}}_{i=1}\frac{z^{2}(A_{i})}{\sigma^{2}_{\text{pred}}(A_{i})}$ (2)
The final calibrated uncertainties are then obtained as $\alpha^{2}\sigma^{2}_{\text{pred}}(A_{i})$ - although in practice we scale the ensemble members around their mean, which simplifies uncertainty propagation. [86] We find that for isotropic chemical shieldings and shielding tensor principal components the calibration factors $\alpha$ are moderate and always close to $1$, indicating that the ShiftML3 ensemble uncertainty estimates are already well calibrated.
We evaluate the uncertainty estimates of the ShiftML3 model on the CSD-test set and plot the predicted uncertainties of 13C shieldings vs the actual prediction errors. In Figure S3 we show parity plots of predicted and calibrated uncertainties of the ShiftML3 model and the actual errors on a hold-out test set.
Figure S3: Predicted uncertainties of 13C isotropic chemical shieldings $\text{std}_{\text{pred}}[\sigma_{\text{iso}}]$ vs the actual prediction errors $\text{z}[\sigma_{\text{iso}}]$ (a) and predicted uncertainties of the corresponding principal components $\text{std}_{\text{pred}}[\sigma_{\text{PAS}}]$ vs the actual component prediction errors $\text{z}[\sigma_{\text{PAS}}]$ (b). The grey lines indicate quantiles of the expected error distribution for a given predicted uncertainty value and are intended to guide the reader’s eye: For example, on the test set, for well calibrated uncertainty estimates there should be 90% of the samples between the two outermost lines as they correspond to quantile lines of 95% and 5% of the log-folded normal distribution.
@@
Only the symmetric part enters directly into the NMR spectrum, whilst the antisymmetric part affects relaxation times. [97] Therefore, in the main part of this manuscript we have focussed on evaluating and utilizing the CSA, $\sigma_{\text{PAS}}$ which are the principal components of $\sigma_{\text{sym}}$. ShiftML3 can predict the full chemical shielding tensor, containing symmetric and antisymmetric parts. In Figure S4 we show parity plots of the symmetric and antisymmetric components $\sigma_{ij,\text{sym}}$ and $\sigma_{ij,\text{antisym}}$ of the 13C and 17O chemical shielding tensors.
@@
In Table S4 we list the prediction accuracies of ShiftML3 on CSD-test separately on antisymmetric and symmetric components of the chemical shielding tensor.
@@
Figure S5: Parity plots of symmetric and antisymmetric chemical shielding tensor components $\sigma_{ij,\text{sym}}$, $\sigma_{ij,\text{antisym}}$ for 13C (a) and 17O (b) - combined parity plots.
@@
Figure S6: Parity plots of principal tensor components $\sigma_{\text{PAS}}$ and symmetric tensor components $\sigma_{ij}$ for 13C (a,b) and 15N (c,d)
In Figure S7 we show parity plots of the principal tensor components resolved for individual principal components $\sigma_{11}$$\sigma_{22}$ and $\sigma_{33}$. In Table S5 we list prediction accuracies of the individual principal components.
@@
Table S5: Predicted accuracy of $\sigma_{\text{PAS}}$ of sML3 on the CSD-test hold-out set, resolved for individual principal components $\sigma_{\text{11}}$, $\sigma_{\text{22}}$ and $\sigma_{\text{33}}$.
Nucleus MAE [ppm] RMSE [ppm] RMSE [%]
@@
For 13C and 15N, the accuracies were evaluated using the benchmark of 36 structures from Ramos et al. [14], with geometries optimized at the PBE-D3(BJ) level of theory [98]. For the evaluation of carbon CSA parameter predictions, we used the benchmark consisting of 19 structures from Hartman et al. [72].
Table S6: RMSE and Linear regression parameters for converting DFT, ShiftML2, and ShiftML3 prediction into chemical shifts for the respective benchmarks.
@@
A direct comparison of the ShiftML3 prediction accuracies with ShiftML2 reveals that it is more accurate for predicting isotropic chemical shieldings across all relevant elements. To the best of our knowledge, no other machine learning model has been trained on the latest iteration of the ShiftML datasets, including thermally distorted structures and additional chemical elements. Unzueta et al. trained GNN models [87] and Liu and coworkers a multiresolution 3D-DenseNet architecture on the CSD-2K dataset [46], an early iteration of the ShiftML training sets. An alternative model construction strategy is the NMRNet from Xu and coworkers, that fine-tuned a transformer model for the chemical shielding prediction in liquids on the CSD-2K set [99]. All three works report evaluation accuracies on the CSD-500 set, a set of DFT relaxed organic crystals, containing at most H,C,N and O. Unfortunately, the number of shared structures between the CSD-test and the CSD-500 test set is too small (about 10 structures) to make direct comparisons between the accuracies of ShiftML3 and other neural network architectures. Shielding computations are highly sensitive to k-space discretization and choice of pseudopotentials [100]. CSD-2K/CSD-500 was computed with other DFT convergence parameters than the ShiftML2/ShiftML3 dataset, severely hampering a direct comparison between all models - we regularly find that shielding values from computations employing different, yet reasonable convergence parameters, differ (in terms of their RMSEs) nearly as much as the ShiftML3 predicted shieldings differ from the CSD-test reference calculations. Instead, we evaluate ShiftML2 and ShiftML3 on 351 structures in our test set that only contain H,C,N,O and are not MD distorted structures (containing 7536 1H, 6546 13C, 665 15N and 1311 17O test environments) and compare their relative improvements in accuracies against the other models, to give some context to the gains in accuracy we achieve with ShiftML3. In Table S7, we compare prediction accuracies of the original ShiftML1.0 model, Unzueta’s ensemble of GNNs, Liu’s MR-3D-DenseNet and compare them with the prediction accuracies of ShiftML2 and ShiftML3 on the H,C,N,O-relaxed sub selection of the CSD-test set.
Table S7: Test set RMSEs (ppm) of the original ShiftML1.0 model (KRR [45]), Unzueta’s GNNs (GNN [87]) and Liu’s MR-3D-DenseNet (DenseNet [46]) and Xu’s NMRNet (NMRNet [99]) on the CSD-500 test set. We compare these values with the prediction accuracies of ShiftML2 (sML2) and ShiftML3 (sML3) on a comparable subset of the CSD-test set. Note that KRR, GNN, DenseNet and NMRNet are evaluated on the CSD-500 test set. As discussed in the text, due to the small overlap of CSD-500 and the CSD-test set, we choose to evaluate sML2 and sML3 on a subset of the CSD-test set containing only H,C,N,O structures that are geometry relaxed. (*) Denotes, that the results of sML2 and sML3 were obtained on that sub selection.